OpenStudy (anonymous):

2 1/2 - 1 3/4 (ESTIMATE)

OpenStudy (anonymous):

is it 3/4?

OpenStudy (anonymous):

if it was 7/4+x=5/2 wouldnt it be 3/4?

OpenStudy (adamaero):

so 5/2 - 7/4 = (10-7)/4 yep

OpenStudy (anonymous):

Thx!